Solve:
5/3x = 4/5 + x

Answer by stanbon(75887)

Solve:
(5/3)x = (4/5) + x
Multiply thru by 15 to get:
25x = 12 + 15x
10x = 12
x = 6/5
x = 1.2
=============
Cheers,
Stan H.